At a certain location close to Earth's surface, we observe a uniform electric field of magnitude 105 N/C directed straight down. What must be the charge (in C) that needs to be placed on a person of mass 75 kg in order to make them lose contact with the ground? Make sure to correctly identify the sign of the charge needed.
Answer #1

Since Weight of person will be always downward, So in order to lose the ground electric force should be in upward direction. Now since electric field is downward, so charge should be negative for electric force to be upward.

Using Force Balance

Fe = Fg

q*E = m*g

q = m*g/E

Using given values:

q = -75 kg*(9.8 m/sec^2)/(10^5 N/C)

q = -7.35*10^-3 C

Please Upvote.